# Extremal behavior of stationary marked point processes

### v2

@@ 4.1 Extremal blocks
where the first equality follows from the refined Campbell’s theorem (2.1). By (3.7),
@@
$\mathbf{P}(X_{\tau,\boldsymbol{i}}\text{ is an extremal cluster})=\mathbf{P}(\textstyle\max_{(t,s)\in X_{\tau,\boldsymbol{i}}}s>a_{\tau}\varepsilon)\\ \leq b_{\tau}^{d}\mathbf{P}(\xi>a_{\tau}\varepsilon)\to 0\quad\text{as}\;\tau\to\infty,$
where we used refined Campbell’s theorem (2.1) in the second step.
@@ 4.2 Space for extremal blocks
Recall that $X$ is a random element of the space $\mathcal{N}_{10}$ so that the blocks $X_{\tau,\boldsymbol{i}}$ can be considered as elements of $\mathbb{N}_{01}$ which consists of simple counting measures on $\mathbb{E}$ with finite values on $\mathbb{R}^{d}\times(\varepsilon,\infty)$ for all $\varepsilon>0$. Recall further that $\mathcal{N}_{01}\subset\mathcal{N}=\mathcal{N}_{11}$ and that, on $\mathcal{N}_{01}$, $\mathcal{B}_{01}$-topology is stronger than the $\mathcal{B}_{11}$-topology. We now define a metric ${\mathsf{m}}$, generating the $\mathcal{B}_{01}$-vague topology on $\mathcal{N}_{01}$.
@@
where the minimum is taken over all permutations $\Pi$ of $\{1,\dots,k\}$. Note that ${\mathsf{m}}_{0}$ is a metric generating the weak (that is, $\mathcal{B}_{00}$-vague) topology on $\mathcal{N}_{00}$, see [25, Proposition 2.3]. While the authors of [25] assume that the ground space is compact, an easy argument justifies the claim for the space $\mathbb{E}=\mathbb{R}^{d}\times(0,\infty)$. Observe also that ${\mathsf{m}}_{0}$ is by construction bounded by $1$ and shift-invariant, that is, ${\mathsf{m}}_{0}(\varphi_{y}\mu,\varphi_{y}\nu)={\mathsf{m}}_{0}(\mu,\nu)$ for all $y\in\mathbb{R}^{d}$ and $\mu,\nu\in\mathcal{N}_{00}$.
@@
where $\mu^{1/u}$ (and similarly $\nu^{1/u}$) is the restriction of $\mu$ on $\mathbb{R}^{d}\times(1/u,\infty)$. One can show that this is indeed a metric on $\mathcal{N}_{01}$ which is bounded by $1$, is shift-invariant, and that it generates the $\mathcal{B}_{01}$-vague topology. The latter claim follows since $\mu_{n}$ converges to $\mu$ in $\mathcal{N}_{01}$ with respect to this topology if and only if $\mu_{n}^{1/u}$ converges weakly to $\mu^{1/u}$ for Lebesgue almost all $u\in(0,\infty)$, see [9, Proposition A2.6.II] and [17].
@@
for all $\mu,\nu\in\mathcal{N}_{01}$. It now follows from [21, Lemma 2.5.1] that $\tilde{{\mathsf{m}}}$ is a pseudo-metric on $\tilde{\mathcal{N}}_{01}$, and that $(\tilde{\mathcal{N}}_{01},\tilde{{\mathsf{m}}})$ is separable since $(\mathcal{N}_{01},{\mathsf{m}})$ is. To show that $\tilde{{\mathsf{m}}}$ is actually a metric, assume that ${\mathsf{m}}(\varphi_{z_{n}}\mu,\nu)\to 0$ for a sequence $(z_{n})$ and conclude that $(z_{n})$ is bounded, so that it admits a convergent subsequence; it is crucial that $\mu$ and $\nu$ have finitely many points in $\mathbb{R}^{d}\times(\varepsilon,\infty)$ for every $\varepsilon$. The second claim follows directly from (4.6). ∎
@@ 4.3 Main result
If Assumptions 4.3, 4.4 and 4.5 below hold, then $\mathbf{P}(Y\in\mathcal{N}_{01})=\mathbf{P}(\Theta\in\mathcal{N}_{01})=1$ and
@@
The second ingredient of the limiting point process in (4.11) is the distribution of $Q$ which can be seen as the asymptotic distribution of a normalized typical cluster of exceedances of $X$. The tail configuration $Y$ can be seen as a cluster of exceedances which contains the uniformly chosen exceedance of $X$ and thus not typical since, compared to the typical cluster, its distribution is biased towards clusters with more exceedances. In fact, the relationship between the tail configuration and the typical cluster of exceedances of $X$ is similar to the relationship between a stationary point process on $\mathbb{R}^{d}$ and its Palm version, and this relationship is discussed in detail in [22] for random fields over $\mathbb{Z}^{d}$.
@@ 4.5 Alternative representations of $\vartheta$ and $Q$
on $\mathcal{N}_{01}$, where $Q$ has distribution (4.13) on $\mathcal{N}_{01}$. Using this fact, one can, as in [22, Proposition 3.9], prove that $Q$ from (4.13) satisfies
@@ 5.1 Small distance to the $k$th nearest neighbor
It is straightforward to see that
$\mathbf{P}(\xi>u)=\mathbf{P}(P(B_{1/u})\geq k)\sim\mathbf{P}(P(B_{1/u})=k)=e^{-C_{d}u^{-d}}\frac{(C_{d}u^{-d})^{k}}{k!}\quad\text{as }\;u\to\infty,$
where $C_{d}$ is the volume of the unit ball in $\mathbb{R}^{d}$. Thus,
@@
Recall that $X=\Psi(P)$ for the scoring function $\psi(s,\mu)=\rho_{k}(s,\mu)^{-1}$, and let $\tilde{P}=\delta_{0}+P$ be a Palm version of $P$. Observe that the conditional distribution of $\tilde{X}$ given $\xi>u$ coincides with the conditional distribution of $\Psi(\tilde{P})$ given $P(B_{1/u})\geq k$. Considering Laplace functionals, it is easy to see that for any nonnegative $f$,
@@
By Proposition 3.2, the tail score at the origin $\eta$ is $\mathrm{Pareto}(dk)$-distributed. Indeed,
@@
$\Theta:=T_{\eta^{-1},\eta}Y=\Big{\{}\big{(}U_{i}/U^{*},\,\rho_{k}(U_{i}/U^{*},\mathcal{Y}/U^{*})^{-1}\big{)},i=0,\dots,k\Big{\}}\,.$
Moreover, it is easy to check that
$\Theta\stackrel{{\scriptstyle d}}{{=}}\Psi(\mathcal{Y}^{*})=\textstyle\sum_{i=0}^{k-1}\delta_{(U_{i},\,\rho_{k}(U_{i},\mathcal{Y}^{*})^{-1})}+\delta_{(U_{k}^{\prime},\,\rho_{k}(U_{k}^{\prime},\mathcal{Y}^{*})^{-1})}\,,$ (5.2)
where $\mathcal{Y}^{*}:=\{0,U_{1},\dots,U_{k-1},U^{\prime}_{k}\}$ with $U^{\prime}_{k}$ uniformly distributed on $\partial B_{1}$ and independent of $U_{1},\dots,U_{k-1}$, which are i.i.d. and uniformly distributed on $B_{1}$.
@@
in all dimensions, where $U_{1}^{\prime\prime}$ is uniform on $\partial B_{1}\cap\{x=(x_{1},\dots,x_{d})\in\mathbb{R}^{d}:x_{1}\geq 0\}$. The value $\vartheta_{1,d}=1/2$ is intuitively obvious since, asymptotically, large values always come in pairs (with exactly the same score). In dimension $d=2$, $\vartheta_{1,2}=1/2$ was obtained in [8, Section 4.2] when analyzing the extremal properties of the inradius of a Poisson–Voronoi tessellation.
@@
In the sixth step we used the refined Campbell’s theorem (2.1), and in the penultimate step we used (4.2) and the fact that $b_{\tau}$ is chosen such that $b_{\tau}/\tau\to 0$. Thus, (4.16) holds, i.e., 4.4 is satisfied.
@@
for all $\boldsymbol{i}\in I_{\tau}$; recall that $\hat{J}_{\tau,\boldsymbol{i}}$ in (4.18) are obtained from the original blocks $J_{\tau,\boldsymbol{i}}$ by trimming the edges by $l_{\tau}$. Thus, since $\widehat{X}_{\tau,\boldsymbol{i}}:=X_{\hat{J}_{\tau,\boldsymbol{i}}}$, the value of $f_{\tau,\boldsymbol{i}}(T_{r(a_{\tau}),a_{\tau}}\widehat{X}_{\tau,\boldsymbol{i}})$ depends only on $P$ restricted to $J_{\tau,\boldsymbol{i}}$, for all $\boldsymbol{i}\in I_{\tau}$. For such $\tau$, since $P$ is a Poisson process and $J_{\tau,\boldsymbol{i}}$’s are disjoint, the left-hand side of (4.19) vanishes, hence, 4.5 holds. ∎
A simple consequence of Proposition 5.3 concerns the behavior od the minimal distance to the $k$-th nearest neighbour in a Poisson configuration of points on an increasing hypercube $[0,\tau]^{d}$. For all $\tau>0$, denote $m_{\rho,\tau}=\min\{\rho_{k}(t,P):t\in P\cap[0,\tau]^{d}\}$. Then, for $a_{\tau}$ as in (5.3), (4.14) implies

### v3

@@ 4.1 Extremal blocks
where the first equality follows from refined Campbell’s theorem (2.1). By (3.7),
@@
$\mathbf{P}(X_{\tau,\boldsymbol{i}}\text{ is an extremal cluster})$ $=\mathbf{P}(\textstyle\max_{(t,s)\in X_{\tau,\boldsymbol{i}}}s>a_{\tau}\varepsilon)$
$\leq b_{\tau}^{d}\mathbf{P}(\xi>a_{\tau}\varepsilon)\to 0\quad\text{as}\;\tau\to\infty,$
where we bound the probability with expectation, then used refined Campbell’s theorem (2.1) and, finally, (4.1).
@@ 4.2 Space for extremal blocks
Recall that $X$ is a random element of the space $\mathcal{N}_{10}$ so that the blocks $X_{\tau,\boldsymbol{i}}$ can be considered as elements of $\mathcal{N}_{01}$ which consists of simple counting measures on $\mathbb{E}$ with finite values on $\mathbb{R}^{d}\times(\varepsilon,\infty)$ for all $\varepsilon>0$. Recall further that $\mathcal{N}_{01}\subset\mathcal{N}=\mathcal{N}_{11}$ and that, on $\mathcal{N}_{01}$, $\mathcal{B}_{01}$-topology is stronger than the $\mathcal{B}_{11}$-topology. We now define a metric ${\mathsf{m}}$, generating the $\mathcal{B}_{01}$-vague topology on $\mathcal{N}_{01}$.
@@
where the minimum is taken over all permutations $\Pi$ of $\{1,\dots,k\}$. Note that ${\mathsf{m}}_{0}$ is a metric generating the weak (that is, $\mathcal{B}_{00}$-vague) topology on $\mathcal{N}_{00}$, see [27, Proposition 2.3]. While the authors of [27] assume that the ground space is compact, an easy argument justifies the claim for the space $\mathbb{E}=\mathbb{R}^{d}\times(0,\infty)$. Observe also that ${\mathsf{m}}_{0}$ is by construction bounded by $1$ and shift-invariant, that is, ${\mathsf{m}}_{0}(\varphi_{y}\mu,\varphi_{y}\nu)={\mathsf{m}}_{0}(\mu,\nu)$ for all $y\in\mathbb{R}^{d}$ and $\mu,\nu\in\mathcal{N}_{00}$.
@@
where $\mu^{1/u}$ (and similarly $\nu^{1/u}$) is the restriction of $\mu$ on $\mathbb{R}^{d}\times(1/u,\infty)$. One can show that this is indeed a metric on $\mathcal{N}_{01}$ which is bounded by $1$, is shift-invariant, and that it generates the $\mathcal{B}_{01}$-vague topology. The latter claim follows since $\mu_{n}$ converges to $\mu$ in $\mathcal{N}_{01}$ with respect to this topology if and only if $\mu_{n}^{1/u}$ converges weakly to $\mu^{1/u}$ for Lebesgue almost all $u\in(0,\infty)$, see [11, Proposition A2.6.II] and [19].
@@
for all $\mu,\nu\in\mathcal{N}_{01}$. It is now easy to show that (4.6) implies that $\tilde{{\mathsf{m}}}$ is a pseudo-metric on $\tilde{\mathcal{N}}_{01}$, and that $(\tilde{\mathcal{N}}_{01},\tilde{{\mathsf{m}}})$ is separable since $(\mathcal{N}_{01},{\mathsf{m}})$ is separable. This follows by a direct application of [23, Lemma 2.5.1]; the only non-trivial step is to show that $\tilde{{\mathsf{m}}}$ satisfies the triangle inequality.
To show that $\tilde{{\mathsf{m}}}$ is actually a metric, assume that $\tilde{{\mathsf{m}}}([\mu],[\nu])=0$ for some $\mu,\nu\in\mathcal{N}_{01}$. By (4.6), there exists a sequence $(z_{n})_{n\in\mathbb{N}}\subseteq\mathbb{R}^{d}$ such that ${\mathsf{m}}(\varphi_{z_{n}}\mu,\nu)\to 0$. If $\nu$ is the null measure, then it follows easily that $\mu$ is also the null measure, so that $[\mu]=[\nu]$. If $\nu$ is not the null measure, for $\varepsilon>0$ small enough, we have that $1\leq\nu(\mathbb{R}^{d}\times(\varepsilon,\infty))<\infty$. Since $\varphi_{z_{n}}$ only translates positions, the sequence $(z_{n})_{n}$ must be bounded. Indeed, otherwise one would have that for infinitely many $n$’s, ${\mathsf{m}}_{0}((\varphi_{z_{n}}\mu)^{1/u},\nu^{1/u})=1$ for $1/u\leq\varepsilon$, and, in particular, ${\mathsf{m}}(\varphi_{z_{n}}\mu,\nu)\geq e^{1/\varepsilon}-1>0$. Thus, there exists a subsequence $(z_{n_{k}})_{k\in\mathbb{N}}$ such that $z_{n_{k}}\to z\in\mathbb{R}^{d}$ as $k\to\infty$. Since $\lim_{k\to\infty}{\mathsf{m}}(\varphi_{z_{n_{k}}}\mu,\varphi_{z}\mu)\to 0$, and since ${\mathsf{m}}$ is a metric, we conclude that $\varphi_{z}\mu=\nu$, i.e. $[\mu]=[\nu]$. ∎
@@ 4.3 Main result
If Assumptions 4.4, 4.5 and 4.6 below hold, then $\mathbf{P}(Y\in\mathcal{N}_{01})=\mathbf{P}(\Theta\in\mathcal{N}_{01})=1$ and
@@
The second ingredient of the limiting point process in (4.11) is the distribution of $Q$ which can be seen as the asymptotic distribution of a normalized typical cluster of exceedances of $X$. In contrast, the tail configuration $Y$ is not typical. Since it contains what can be intuitively understood as a uniformly selected exceedance in $X$, the distribution of $Y$ is biased towards clusters with more exceedances. In fact, the relationship between the tail configuration and the typical cluster of exceedances of $X$ is similar to the relationship between a stationary point process on $\mathbb{R}^{d}$ and its Palm version, and this relationship is discussed in detail in [24] for random fields over $\mathbb{Z}^{d}$.
Figure 1: Simulated exceedances of scores over a fixed threshold and an illustration of the block limit of $N_{\tau}$ in Theorem 4.2 for the case with and without clustering. The model is taken from Example 1.1 with $p=1/2$ and $p=0$.
Example 4.3 (continuation of Example 1.1).
Recall that for the stationary point process $X$ of our initial example, we found the tail configuration in the form $Y=\delta_{(0,\eta)}+I\delta_{(h_{0},\eta)}$, with the distribution of $I,\,h_{0}$ and $\eta$ described above. By (3.9) the spectral tail configuration has the form $\Theta=\delta_{(0,1)}+I\delta_{(h_{0},1)}$. For $p=1/2$ we have $\mathbf{P}(I=1)=2/3$, therefore,
$\vartheta=\mathbf{P}(A^{\mathrm{fm}}(\Theta)=0)=\mathbf{P}(I=0)+\mathbf{P}(I=1)\mathbf{P}(\min(0,h_{0})=0)=2/3.$
Finally, it is not difficult to show that $Q=\delta_{(0,1)}+\varepsilon\delta_{(h^{+}_{0},1)}$ where $h^{+}_{0}$ has the distribution of $h_{0}$ restricted to the points larger than 0 in the lexicographical order on $\mathbb{R}^{d}$ and $\varepsilon$ is an independent Bernoulli random variable with $\mathbf{P}(\varepsilon=1)=1/2$.
@@
Assumption 4.5 concerns the maximum score of $\tilde{X}$ in the annulus $C_{\tau,u}=B_{b_{\tau}c}\setminus B_{r(a_{\tau}\varepsilon)u}$, which, for a sufficiently large $u$, is far away from the origin and still relatively small compared to the size of $[0,\tau]^{d}$. The assumption simply states that if we condition on a large score at the origin, we are unlikely to see another large value in such an area. Roughly speaking, it prevents clustering of large scores beyond a certain distance. For a further illustration of this condition see Example 4.7.
@@
Example 4.7 (continuation of Example 1.1).
Recall that $X=\sum\delta_{(t,\zeta_{t})}+\varepsilon_{t}\delta_{(t+h_{t},\zeta_{t})}$ where i.i.d. points $(h_{t},\varepsilon_{t},\zeta_{t})$ have independent components with the first having bounded support, i.e. $\mathbf{P}(\|h\|\leq H)=1$ for some constant $H\geq 0$. If we set $r\equiv 1$, Assumption 4.4 holds for any $b_{\tau}\to\infty$ and $b_{\tau}=o(\tau)$. Assumption 4.5 is also easily verified in this context: simply recall that $\tilde{X}=X+C_{0}$ with $C_{0}\stackrel{{\scriptstyle d}}{{=}}\delta_{(0,\xi)}+I\delta_{(h_{0},\xi)}$. Therefore, for any $u>H$, there is no point of $C_{0}$ in the annulus $C_{\tau,u}=B_{b_{\tau}c}\setminus B_{u}$ and thus
$\mathbf{P}(M(\tilde{X}_{C_{\tau,u}})>a_{\tau}\delta\mid\xi>a_{\tau}\varepsilon)=\mathbf{P}(M(X_{C_{\tau,u}})>a_{\tau}\delta)\leq\mathrm{const}\cdot b_{\tau}^{d}\mathbf{P}(\zeta>a_{\tau}\delta)\to 0$
as $\tau\to\infty$. Similarly, because any cluster fits in a ball of radius $H$, Assumption 4.6 holds immediately, if we trim the block with $l_{\tau}=H$.
@@ 4.5 Alternative representations of $\vartheta$ and $Q$
on $\mathcal{N}_{01}$, where $Q$ has distribution (4.13) on $\mathcal{N}_{01}$. Using this fact, one can, as in [24, Proposition 3.9], prove that $Q$ from (4.13) satisfies
@@ 5.1 Small distance to the $k$th nearest neighbor
Below we consider the situation when each point of the stationary Poisson process is equipped with the score, being the reciprocal of the distance to its nearest neighbour. Then large scores identify points with small distances to nearest neighbours and the tail configuration describes the positions of points in a cluster of points which all are located near to each other. Note that large distances to the nearest neighbour recently studied in [8] indicate isolated points which do not form clusters.
@@
Since the random variable $P(B_{1/u})$ has Poisson distribution with mean $C_{d}u^{-d}$, where $C_{d}$ is the volume of the unit ball in $\mathbb{R}^{d}$, and this mean goes to zero as $u\to\infty$, it is straightforward to see that
$\mathbf{P}(\xi>u)=\mathbf{P}(P(B_{1/u})\geq k)\sim\mathbf{P}(P(B_{1/u})=k)=e^{-C_{d}u^{-d}}\frac{(C_{d}u^{-d})^{k}}{k!}\quad\text{as }\;u\to\infty.$
Thus,
@@
Recall that $X=\Psi(P)$ for the scoring function $\psi(s,\mu)=\rho_{k}(s,\mu)^{-1}$, and let $\tilde{P}=\delta_{0}+P$ be a Palm version of $P$. Observe that the conditional distribution of $\tilde{X}$ given $\xi>u$ coincides with the conditional distribution of $\Psi(\tilde{P})$ given $P(B_{1/u})\geq k$. In view of this and taking into account that $\mathbf{P}(P(B_{1/u})\geq k)\sim\mathbf{P}(P(B_{1/u})=k)$, it is easy to see that for any nonnegative $f$ the Laplace functional of $\tilde{X}$ conditional on $\{\xi>u\}$ satisfies
@@
By Proposition 3.3, the tail score at the origin $\eta$ is $\mathrm{Pareto}(dk)$-distributed which is easily checked since
@@
$\Theta:=T_{\eta^{-1},\eta}Y$ =$=\Big{\{}\big{(}U_{i}/U^{*},\,\rho_{k}(U_{i}/U^{*},\mathcal{Y}/U^{*})^{-1}\big{)},i=0,\dots,k\Big{\}}$
$=\Psi(\{U_{0}/U^{*},\dots,U_{k}/U^{*}\})\,.$
A direct calculation shows that the random set $\{U_{0}/U^{*},\dots,U_{k}/U^{*}\}$ has the same distribution as $\mathcal{Y}^{*}:=\{U_{0},U_{1},\dots,U_{k-1},U^{\prime}_{k}\}$, where $U^{\prime}_{k}$ is uniformly distributed on $\partial B_{1}$ and independent of $U_{0},U_{1},\dots,U_{k-1}$. In particular,
$\Theta\stackrel{{\scriptstyle d}}{{=}}\Psi(\mathcal{Y}^{*})=\textstyle\sum_{i=0}^{k-1}\delta_{(U_{i},\,\rho_{k}(U_{i},\mathcal{Y}^{*})^{-1})}+\delta_{(U_{k}^{\prime},\,\rho_{k}(U_{k}^{\prime},\mathcal{Y}^{*})^{-1})}\,.$ (5.2)
@@
in all dimensions, where $U_{1}^{\prime\prime}$ is uniform on $\partial B_{1}\cap\{x=(x_{1},\dots,x_{d})\in\mathbb{R}^{d}:x_{1}\geq 0\}$. The value $\vartheta_{1,d}=1/2$ is intuitively obvious since, asymptotically, large values always come in pairs (with exactly the same score). In dimension $d=2$, $\vartheta_{1,2}=1/2$ was obtained in [9, Section 4.2] when analyzing the extremal properties of the inradius of a Poisson–Voronoi tessellation.
@@
In the sixth step we used the refined Campbell’s theorem (2.1), and in the penultimate step we used (4.2) and the fact that $b_{\tau}$ is chosen such that $b_{\tau}/\tau\to 0$. Thus, (4.16) holds, i.e., 4.5 is satisfied.
@@
for all $\boldsymbol{i}\in I_{\tau}$; recall that $\hat{J}_{\tau,\boldsymbol{i}}$ in (4.18) are obtained from the original blocks $J_{\tau,\boldsymbol{i}}$ by trimming the edges by $l_{\tau}$. Thus, since $\widehat{X}_{\tau,\boldsymbol{i}}:=X_{\hat{J}_{\tau,\boldsymbol{i}}}$, the value of $f_{\tau,\boldsymbol{i}}(T_{r(a_{\tau}),a_{\tau}}\widehat{X}_{\tau,\boldsymbol{i}})$ depends only on $P$ restricted to $J_{\tau,\boldsymbol{i}}$, for all $\boldsymbol{i}\in I_{\tau}$. For such $\tau$, since $P$ is a Poisson process and $J_{\tau,\boldsymbol{i}}$’s are disjoint, the left-hand side of (4.19) vanishes, hence, 4.6 holds. ∎
A simple consequence of Proposition 5.3 concerns the behavior of the minimal distance to the $k$-th nearest neighbour in a Poisson configuration of points on an increasing hypercube $[0,\tau]^{d}$. For all $\tau>0$, denote $m_{\rho,\tau}=\min\{\rho_{k}(t,P):t\in P\cap[0,\tau]^{d}\}$. Then, for $a_{\tau}$ as in (5.3), (4.14) implies